# Whitehead’s problem and condensed mathematics

### v3

@@ 11. A different vantage point
To see the fifth equality, apply Lemma 3.2; for the third and fourth, respectively, apply [18, Prop. 3.5 and Prop. 3.2]. The first equality follows from [26, p. 13 (iii) (cf. p. 25)]. This leaves us only with the second. Here the idea, evidently, is to apply the Hom-tensor adjunction at the level of $\mathrm{Ext}^{1}_{\mathsf{CondAb}}$, and we may do so by the naturality of the adjunction at the $\mathrm{Hom}_{\mathsf{CondAb}}$-level, the flatness of $\mathbb{Z}[\underline{S}]$ [26, p. 13], and the fact, noted already, that $\underline{\bigoplus_{I}\mathbb{Z}}=\bigoplus_{I}\underline{\mathbb{Z}}$ for any set $I$. More particularly, letting $\underline{\mathcal{P}}$, much as in the proof of Lemma 3.2, denote the condensation of a free resolution of $\underline{A}$, we have:
@@
and this chain of equalities, if sound, will complete the argument. Let us check the details. Observe first that $\underline{\mathcal{P}}$ is a projective resolution of $\underline{A}$, by [18, Prop. 2.18] and Proposition 12.3 below; this shows the third equality. The second is the aforementioned adjunction. For the first, we have noted already that $\mathbb{Z}[\underline{S}]$ is flat, so the question reduces to that of whether the terms of $\underline{\mathcal{P}}\otimes\mathbb{Z}[\underline{S}]$ are projective. For this observe simply that $\big{(}\underline{\bigoplus_{I}\mathbb{Z}}\big{)}\otimes\mathbb{Z}[\underline{S}]=\big{(}\bigoplus_{I}\underline{\mathbb{Z}}\big{)}\otimes\mathbb{Z}[\underline{S}]=\bigoplus_{I}\big{(}\underline{\mathbb{Z}}\otimes\mathbb{Z}[\underline{S}]\big{)}=\bigoplus_{I}\mathbb{Z}[\underline{S}]$ for any set $I$; that such sums are projective follows from Proposition 12.3 below. ∎
@@
is independent of the $\mathsf{ZFC}$ axioms. Clausen and Scholze’s result, in contrast, can now be seen to follow from the fact that the implication
$\big{[}\,\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,C(S,\mathbb{Z}))=0\text{ for all }S\in\mathsf{ED}\,\big{]}\;\Rightarrow\;\text{the abelian group }A\text{ is free}$
is indeed a $\mathsf{ZFC}$ theorem. This implication is, moreover, straightforward to prove, since by a result of Nöbeling ([22], see also [10, Proposition XI.4.4]), generalizing work of Specker [29], the group $C(S,\mathbb{Z})$ is itself a free abelian group, one isomorphic to the free abelian group on the topological weight of $S$.
@@
It follows that $\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,\mathbb{Z}^{(\kappa)})\neq 0$. Therefore, Lemma 11.1 immediately yields the following corollary, providing an alternative proof of clauses (1) and (3) of Theorem 1.3.
@@
We first note that, by results in [11], it is consistent with the axioms of $\mathsf{ZFC}$ that, for every cardinal $\kappa$, there is a nonfree abelian group $A$ such that $\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,B)=0$ for every abelian group $B$ of cardinality ${\leq}\kappa$. Thus, Lemma 11.1 yields the following corollary, establishing clause (2) of Theorem 1.4 and indicating that, consistently, the positive answer to the interpretation of Whitehead’s problem in $\mathsf{CondAb}$ established by Theorem 1.2 fails in $\mathsf{CondAb}_{\kappa}$ for every uncountable cardinal $\kappa$.
@@
It is consistent with the axioms of $\mathsf{ZFC}$ that, for every cardinal $\kappa$, there is a nonfree abelian group $A$ such that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)=0$ for all $S\in\mathsf{ED}_{\kappa}$. In particular, it is consistent that, for every uncountable cardinal $\kappa$, there is a nonfree abelian group $A$ such that $\underline{A}$ is Whitehead in $\mathsf{CondAb}_{\kappa}$.
@@
Suppose that $\underline{A}$ is Whitehead in $\mathsf{CondAb}_{\kappa}$. In particular, for for every $S\in\mathsf{ED}_{\kappa}$, we have $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)=0$. By Lemma 11.1, it follows that, for every cardinal $\lambda<\kappa$, we have $\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,\mathbb{Z}^{(\lambda)})=0$. By the observations in the paragraph immediately preceding Corollary 11.2, it follows that $A$ is $\kappa$-free, i.e., every subgroup of $A$ of cardinality ${<}\kappa$ is free. By [10, Corollary II.3.11], since $\kappa$ is strongly compact, $A$ is free, as desired. [^5] As [10, Corollary II.3.11] makes clear, we could have merely assumed that $\kappa$ is only $L_{\omega_{1}\omega}$-compact, a weaker but lesser-known hypothesis than the strong compactness of $\kappa$.
@@
In the proof of Corollary 11.4, the crucial property of the cardinal $\kappa$ was the fact that every $\kappa$-free group is free. This is true if $\kappa$ is a strongly compact cardinal, but we remark that, consistently, this can also be true for smaller values of $\kappa$. In particular, it is shown in [16, §4] that, assuming the consistency of the existence of infinitely many supercompact cardinals, there exists a model of $\sf ZFC$ in which this condition holds for the first cardinal fixed point of the $\aleph$ function, i.e., the least $\kappa$ such that $\kappa=\aleph_{\kappa}$.
@@ 12. Locally compact and solid abelian groups
In this section, we show that Theorem 1.2 extends from the category of (discrete) abelian groups to the category of locally compact abelian groups; here it is useful to recall that the projective objects of the former category are exactly the free ones. In particular, we prove the following theorem.
@@
By the structure theorem for locally compact abelian groups (see [26, Thm. 4.1(i)]), any locally compact abelian group $A$ is of the form $\mathbb{R}^{n}\times B$ for some nonnegative integer $n$, where $B$ is an extension of a compact group by a discrete one, or where, in other words, there exists a short exact sequence
@@
with $C$ compact and $D$ discrete whose image in $\mathsf{CondAb}$ is exact as well (see [18, Props. 4.9, 2.18]). Since $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{\mathbb{R}},\underline{\mathbb{Z}})=0$ [26, p. 25], we have
@@
(The rightmost 00 follows from [26, p. 27] (e.g., from the comment following Cor. 4.9); the leftmost 00 follows from [26, Prop. 4.2].) Hence
@@
By Theorem 1.2, item (1) holds if and only if $D=\bigoplus_{I}\mathbb{Z}$ for some $I$. Also, the argument of [18, Prop. 4.11] together with [26, Prop. 4.3(ii)] shows that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{R}})=0$; this together with the fact that $\underline{\mathrm{Hom}}(\underline{C},\underline{\mathbb{R}})=\underline{\mathrm{Hom}(C,\mathbb{R})}=0$ ([26, Prop. 4.2]) implies that the map $\underline{\mathrm{Hom}}(\underline{C},\underline{\mathbb{R}/\mathbb{Z}})\to\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{Z}})$ appearing in the long exact sequence induced by applying $\underline{\mathrm{Hom}}(\underline{C},-)$ to $\underline{\mathbb{Z}}\to\underline{\mathbb{R}}\to\underline{\mathbb{R}/\mathbb{Z}}$ is an isomorphism. Again applying [26, Prop. 4.2], we have $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{Z}})=\underline{\mathrm{Hom}(C,\mathbb{R}/\mathbb{Z})}=\underline{\mathbb{D}(C)}$, the condensed image of the Pontryagin dual group of $C$; in particular, item (2) holds if and only if $C=0$. Summing up, $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})=0$ if and only if $A$ is of the form $\mathbb{R}^{n}\times D$ for some $D=\bigoplus_{I}\mathbb{Z}$. As noted, by [21, Prop. 3.3], such $A$ are precisely the projective objects of $\mathsf{LCA}$. ∎
@@
By the lemma, any section $s:\underline{\mathbb{R}}\to\bigoplus_{I}\mathbb{Z}[\underline{S_{i}}]$ of an $r$ as above would evaluate over a point $\ast$ as an injection $s(*):\mathbb{R}\to\bigoplus_{I}\bigoplus_{S_{i}}\mathbb{Z}$, but any nonzero element in the image of $s(*)$ would need to be divisible by every positive $n$, a contradiction. The following, therefore, will complete the proof of Corollary 12.2.
@@
Recall also that the sheafification $\mathcal{F}^{\sharp}$ of a presheaf $\mathcal{F}$ on a site is given by the twofold iteration of the $+$ operation, where
@@
Theorem 12.1 opens onto broader questions about alternative interpretations of Whitehead’s problem in the condensed setting. These may be thought of as defined by two main parameters: the subscript $s$ of the $\underline{\mathrm{Ext}}^{1}$ term, and the domain $d$ of $\underline{\mathrm{Ext}}^{1}_{s}(\cdot,\underline{\mathbb{Z}})$. After all, in retrospect, it is not altogether surprising that ($s=\mathsf{CondAb}$)-valued invariants $\underline{\mathrm{Ext}}^{1}_{s}$ should be strong enough to settle the problem over the comparatively weaker domains of $d=\mathsf{Ab}$ or even $d=\mathsf{LCA}$ (we hope to have shown that it is altogether interesting nevertheless). A natural next question, then, is how much further this $d$ may be pushed. By Corollary 12.2, of course, Whitehead’s problem has a ZFC solution when $d=\mathsf{CondAb}$ as well, but it is an unsatisfying one, deriving as it does from the arguably spurious case of $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{\mathbb{R}},\underline{\mathbb{Z}})=0$. Against this background, particularly interesting is the problem when $d$ equals the subcategory $\mathsf{Solid}$ of $\mathsf{CondAb}$ mentioned briefly in our introduction. Forming, as it does, the paradigmatic example of an analytic ring, this is one of $\mathsf{CondAb}$’s most critical subcategories (see [26, §5–7] and [25, §2]); moreover, as noted, it was this category’s features which Clausen and Scholze leveraged in their original proof of Theorem 1.2. The solid setting, being “very much a ‘nonarchimedean’ notion” [26, footnote 10 on p. 32], is also one in which $\underline{\mathbb{R}}$ does not appear. [^6] More precisely, the inclusion $\mathsf{Solid}\to\mathsf{CondAb}$ possesses a left adjoint solidification functor notated $M\mapsto M^{\blacksquare}$, and the solidification $\underline{\mathbb{R}}^{\blacksquare}$ of $\underline{\mathbb{R}}$ is 00 [26, Corollary 6.1(iii)]. Within it, lastly, $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}$ and $\underline{\mathrm{Ext}}^{1}_{\mathsf{Solid}}$ coincide [26, Lemma 5.9], so that this instance of the problem recovers the $s=d$ symmetry of the classical one:
@@
where $\mathbf{A}[H]$ denotes for any abelian group $H$ the inverse system indexed by functions $f:\mathbb{N}\to\mathbb{N}$, with terms $A_{f}:=\bigoplus_{i\in\mathbb{N}}H^{(f(i))}$ and natural projection maps $A_{g}\to A_{f}$ for any $f$ which is pointwise dominated by $g$. See [5] for details of this conversion; the immediate point, though, is that these derived limits have been the subject of much recent study (see [4, 3, 33]), from which the consistency with the ZFC axioms of the statement that
@@
is also consistent with the ZFC axioms [1, Thm. 3]. By the following argument, though, $A$ witnesses a negative answer to Question 12.5 in this scenario as well.
@@
defined as follows: fix $g:\mathbb{N}\to\mathbb{N}$ and an element $a$ of the $g^{\mathrm{th}}$ summand of the domain of $t$; clearly it will suffice to describe the effect of $t$ on such $a$. Here it is convenient to adopt the (elsewhere innocuous) convention that 00 is not a natural number: the first $g(1)$ many entries in $a$ are then mapped to the first $g(1)$ summands of the first factor of $\prod_{\mathbb{N}}\bigoplus_{\mathbb{N}}\mathbb{Z}$, the next $g(2)$ many entries to the first $g(2)$ many summands of the second factor of $\prod_{\mathbb{N}}\bigoplus_{\mathbb{N}}\mathbb{Z}$, and so on. Equipping $\mathbb{Z}$ with the discrete topology and letting products and sums induce the product and coproduct topologies, respectively, we may view $t$ as a continuous mapping of topological abelian groups. Consider the condensed image $\underline{t}$ of $t$: observe that its codomain is $A$, its domain is solid, and that it remains an epimorphism (cf. [18, Prop. 2.18]). Thus, to show that $A$ is not projective in $\mathsf{Solid}$, it will suffice to show that $\underline{t}$ does not split, and for this it will suffice to show that the mapping $\underline{t}(*)=t$ of discrete abelian groups described above possesses no right-inverse $s$. By [10, Cor. III.1.9], though, for any
@@
Even this doesn’t altogether settle the question, though, since it is also consistent with the ZFC axioms (and follows from the continuum hypothesis) that $\underline{\mathrm{Ext}}^{1}(A,\underline{\mathbb{Z}})\neq 0$ [17, Thm. 2]. For a model-independent negative answer to Question 12.5, take the solid abelian group $B=\underline{\mathbb{Z}_{p}}/\underline{\mathbb{Z}[1/q]}$ for any distinct prime numbers $p$ and $q$. As noted in Clausen’s Lecture 11 of [8],
@@
(see [19, Rmk. 2, p. 2] for the first equality), and for its denominator
@@
Hence, applying $R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(\,\cdot\,,\underline{\mathbb{Z}})$ to the short exact sequence

### v5

@@ 11. A different vantage point
To see the fifth equality, apply Lemma 3.2; for the third and fourth, respectively, apply [19, Prop. 3.5 and Prop. 3.2]. The first equality follows from [28, p. 13 (iii) (cf. p. 25)]. This leaves us only with the second. Here the idea, evidently, is to apply the Hom-tensor adjunction at the level of $\mathrm{Ext}^{1}_{\mathsf{CondAb}}$, and we may do so by the naturality of the adjunction at the $\mathrm{Hom}_{\mathsf{CondAb}}$-level, the flatness of $\mathbb{Z}[\underline{S}]$ [28, p. 13], and the fact, noted already, that $\underline{\bigoplus_{I}\mathbb{Z}}=\bigoplus_{I}\underline{\mathbb{Z}}$ for any set $I$. More particularly, letting $\underline{\mathcal{P}}$, much as in the proof of Lemma 3.2, denote the condensation of a free resolution of $\underline{A}$, we have:
@@
and this chain of equalities, if sound, will complete the argument. Let us check the details. Observe first that $\underline{\mathcal{P}}$ is a projective resolution of $\underline{A}$, by [19, Prop. 2.18] and Proposition 12.3 below; this shows the third equality. The second is the aforementioned adjunction. For the first, we have noted already that $\mathbb{Z}[\underline{S}]$ is flat, so the question reduces to that of whether the terms of $\underline{\mathcal{P}}\otimes\mathbb{Z}[\underline{S}]$ are projective. For this observe simply that $\big(\underline{\bigoplus_{I}\mathbb{Z}}\big)\otimes\mathbb{Z}[\underline{S}]=\big(\bigoplus_{I}\underline{\mathbb{Z}}\big)\otimes\mathbb{Z}[\underline{S}]=\bigoplus_{I}\big(\underline{\mathbb{Z}}\otimes\mathbb{Z}[\underline{S}]\big)=\bigoplus_{I}\mathbb{Z}[\underline{S}]$ for any set $I$; that such sums are projective follows from Proposition 12.3 below. ∎
@@
is independent of the $\sf ZFC$ axioms. Clausen and Scholze’s result, in contrast, can now be seen to follow from the fact that the implication
[…]
is indeed a $\sf ZFC$ theorem. This implication is, moreover, straightforward to prove, since by a result of Nöbeling ([23], see also [10, Proposition XI.4.4]), generalizing work of Specker [31], the group $C(S,\mathbb{Z})$ is itself a free abelian group, one isomorphic to the free abelian group on the topological weight of $S$.
@@
It follows that $\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,\mathbb{Z}^{(\kappa)})\neq 0$. Therefore, Lemma 11.1 immediately yields the following corollary, providing an alternative proof of clause (2) of Theorem 1.3.
@@
We first note that, by results in [11], it is consistent with the axioms of $\sf ZFC$ that, for every cardinal $\kappa$, there is a nonfree abelian group $A$ such that $\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,B)=0$ for every abelian group $B$ of cardinality ${\leq}\kappa$. Thus, Lemma 11.1 yields the following corollary, establishing clause (2) of Theorem 1.4 and indicating that, consistently, the positive answer to the interpretation of Whitehead’s problem in $\mathsf{CondAb}$ established by Theorem 1.2 fails in $\mathsf{CondAb}_{\kappa}$ for every uncountable cardinal $\kappa$.
@@
It is consistent with the axioms of $\sf ZFC$ that, for every cardinal $\kappa$, there is a nonfree abelian group $A$ such that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)=0$ for all $S\in\mathsf{ED}_{\kappa}$. In particular, it is consistent that, for every uncountable cardinal $\kappa$, there is a nonfree abelian group $A$ such that $\underline{A}$ is Whitehead in $\mathsf{CondAb}_{\kappa}$.
@@
Suppose that $\underline{A}$ is Whitehead in $\mathsf{CondAb}_{\kappa}$. In particular, for for every $S\in\mathsf{ED}_{\kappa}$, we have $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)=0$. By Lemma 11.1 and the aforementioned result of Nöbeling indicating that, for every $S\in\mathsf{ED}$, $C(S,\mathbb{Z})$ is the free abelian group on the topological weight of $S$, it follows that, for every cardinal $\lambda<\kappa$, we have $\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,\mathbb{Z}^{(\lambda)})=0$. By the observations in the paragraph immediately preceding Corollary 11.2, it follows that $A$ is $\kappa$-free, i.e., every subgroup of $A$ of cardinality $\kappa$ is free. By [10, Corollary II.3.11], since $\kappa$ is strongly compact, $A$ is free, as desired. [^6] As [10, Corollary II.3.11] makes clear, we could have merely assumed that $\kappa$ is only $L_{\omega_{1}\omega}$-compact, a weaker but lesser-known hypothesis than the strong compactness of $\kappa$.
@@
In the proof of Corollary 11.4, the crucial property of the cardinal $\kappa$ was the fact that every $\kappa$-free group is free. This is true if $\kappa$ is a strongly compact cardinal, but we remark that, consistently, this can also be true for smaller values of $\kappa$. In particular, it is shown in [17, §4] that, assuming the consistency of the existence of infinitely many supercompact cardinals, there exists a model of $\sf ZFC$ in which this condition holds for the first cardinal fixed point of the $\aleph$ function, i.e., the least $\kappa$ such that $\kappa=\aleph_{\kappa}$.
@@ 12. Locally compact and solid abelian groups
In this section, we prove Theorem 1.5, which we restate here for the reader’s convenience.
@@
By the structure theorem for locally compact abelian groups (see [28, Thm. 4.1(i)]), any locally compact abelian group $A$ is of the form $\mathbb{R}^{n}\times B$ for some nonnegative integer $n$, where $B$ is an extension of a compact group by a discrete one, or where, in other words, there exists a short exact sequence
@@
with $C$ compact and $D$ discrete whose image in $\mathsf{CondAb}$ is exact as well (see [19, Props. 4.9, 2.18]). Since $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{\mathbb{R}},\underline{\mathbb{Z}})=0$ [28, p. 25], we have
@@
(The rightmost 0 follows from [28, p. 27] (e.g., from the comment following Cor. 4.9); the leftmost 0 follows from [28, Prop. 4.2].) Hence
@@
By Theorem 1.2, item (1) holds if and only if $D=\bigoplus_{I}\mathbb{Z}$ for some $I$. Also, the argument of [19, Prop. 4.11] together with [28, Prop. 4.3(ii)] shows that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{R}})=0$; this together with the fact that $\underline{\mathrm{Hom}}(\underline{C},\underline{\mathbb{R}})=\underline{\mathrm{Hom}(C,\mathbb{R})}=0$ ([28, Prop. 4.2]) implies that the map $\underline{\mathrm{Hom}}(\underline{C},\underline{\mathbb{R}/\mathbb{Z}})\to\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{Z}})$ appearing in the long exact sequence induced by applying $\underline{\mathrm{Hom}}(\underline{C},-)$ to $\underline{\mathbb{Z}}\to\underline{\mathbb{R}}\to\underline$$\mathbb{R}$$\mathbb{Z}$} is an isomorphism. Again applying [28, Prop. 4.2], we have $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{Z}})=\underline{\mathrm{Hom}(C,\mathbb{R}/\mathbb{Z})}=\underline{\mathbb{D}(C)}$, the condensed image of the Pontryagin dual group of $C$; in particular, item (2) holds if and only if $C=0$. Summing up, $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})=0$ if and only if $A$ is of the form $\mathbb{R}^{n}\times D$ for some $D=\bigoplus_{I}\mathbb{Z}$. As noted, by [22, Prop. 3.3], such $A$ are precisely the projective objects of $\mathsf{LCA}$. ∎
@@
By the lemma, any section $s:\underline{\mathbb{R}}\to\bigoplus_{I}\mathbb{Z}[\underline{S_{i}}]$ of an $r$ as above would evaluate over a point $\ast$ as an injection $s(*):\mathbb{R}\to\bigoplus_{I}\bigoplus_{S_{i}}\mathbb{Z}$, but any nonzero element in the image of s()s(*) would need to be divisible by every positive $n$, a contradiction. The following, therefore, will complete the proof of Corollary 12.2.
@@
Recall also that the sheafification $\mathcal{F}^{\sharp}$ of a presheaf $\mathcal{F}$ on a site is given by the twofold iteration of the ++ operation, where
@@
Theorem 12.1 opens onto broader questions about alternative interpretations of Whitehead’s problem in the condensed setting. These may be thought of as defined by two main parameters: the subscript $s$ of the $\underline{\mathrm{Ext}}^{1}$ term, and the domain $d$ of $\underline{\mathrm{Ext}}^{1}_{s}(\cdot,\underline{\mathbb{Z}})$. After all, in retrospect, it is not altogether surprising that ($s=\mathsf{CondAb}$)-valued invariants $\underline{\mathrm{Ext}}^{1}_{s}$ should be strong enough to settle the problem over the comparatively weaker domains of $d=\mathsf{Ab}$ or even $d=\mathsf{LCA}$ (we hope to have shown that it is altogether interesting nevertheless). A natural next question, then, is how much further this $d$ may be pushed. By Corollary 12.2, of course, Whitehead’s problem has a $\sf ZFC$ solution when $d=\mathsf{CondAb}$ as well, but it is an unsatisfying one, deriving as it does from the arguably spurious case of $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{\mathbb{R}},\underline{\mathbb{Z}})=0$. Against this background, particularly interesting is the problem when $d$ equals the subcategory $\mathsf{Solid}$ of $\mathsf{CondAb}$ mentioned briefly in our introduction. Forming, as it does, the paradigmatic example of an analytic ring, this is one of $\mathsf{CondAb}$’s most critical subcategories (see [28, §5–7] and [27, §2]); moreover, as noted, it was this category’s features which Clausen and Scholze leveraged in their original proof of Theorem 1.2. The solid setting, being “very much a ‘nonarchimedean’ notion” [28, footnote 10 on p. 32], is also one in which $\underline{\mathbb{R}}$ does not appear. [^7] More precisely, the inclusion $\mathsf{Solid}\to\mathsf{CondAb}$ possesses a left adjoint solidification functor notated $M\mapsto M^{\blacksquare}$, and the solidification $\underline{\mathbb{R}}^{\blacksquare}$ of $\underline{\mathbb{R}}$ is 0 [28, Corollary 6.1(iii)]. Within it, lastly, $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}$ and $\underline{\mathrm{Ext}}^{1}_{\mathsf{Solid}}$ coincide [28, Lemma 5.9], so that this instance of the problem recovers the $s=d$ symmetry of the classical one:
@@
where $\mathbf{A}[H]$ denotes for any abelian group $H$ the inverse system indexed by functions $f:\mathbb{N}\to\mathbb{N}$, with terms $A_{f}:=\bigoplus_{i\in\mathbb{N}}H^{(f(i))}$ and natural projection maps $A_{g}\to A_{f}$ for any $f$ which is pointwise dominated by $g$. See [5] for details of this conversion; the immediate point, though, is that these derived limits have been the subject of much recent study (see [4, 3, 35]), from which the consistency with the $\sf ZFC$ axioms of the statement that
@@
is also consistent with the $\sf ZFC$ axioms [1, Thm. 3]. By the following argument, though, $A$ witnesses a negative answer to Question 12.5 in this scenario as well.
@@
defined as follows: fix $g:\mathbb{N}\to\mathbb{N}$ and an element $a$ of the $g^{\mathrm{th}}$ summand of the domain of $t$; clearly it will suffice to describe the effect of $t$ on such $a$. Here it is convenient to adopt the (elsewhere innocuous) convention that 0 is not a natural number: the first $g(1)$ many entries in $a$ are then mapped to the first $g(1)$ summands of the first factor of $\prod_{\mathbb{N}}\bigoplus_{\mathbb{N}}\mathbb{Z}$, the next $g(2)$ many entries to the first $g(2)$ many summands of the second factor of $\prod_{\mathbb{N}}\bigoplus_{\mathbb{N}}\mathbb{Z}$, and so on. Equipping $\mathbb{Z}$ with the discrete topology and letting products and sums induce the product and coproduct topologies, respectively, we may view $t$ as a continuous mapping of topological abelian groups. Consider the condensed image $\underline{t}$ of $t$: observe that its codomain is $A$, its domain is solid, and that it remains an epimorphism (cf. [19, Prop. 2.18]). Thus, to show that $A$ is not projective in $\mathsf{Solid}$, it will suffice to show that $\underline{t}$ does not split, and for this it will suffice to show that the mapping $\underline{t}(*)=t$ of discrete abelian groups described above possesses no right-inverse $s$. By [10, Cor. III.1.9], though, for any
@@
Even this doesn’t altogether settle the question, though, since it is also consistent with the $\sf ZFC$ axioms (and follows from the continuum hypothesis) that $\underline{\mathrm{Ext}}^{1}(A,\underline{\mathbb{Z}})\neq 0$ [18, Thm. 2]. For a model-independent negative answer to Question 12.5, take the solid abelian group $B=\underline{\mathbb{Z}_{p}}/\underline{\mathbb{Z}[1/q]}$ for any distinct prime numbers $p$ and $q$. As noted in Clausen’s Lecture 11 of [8],
@@
(see [20, Rmk. 2, p. 2] for the first equality), and for its denominator
@@
[…]